OpenStudy (anonymous):

Solve: −2(x − 1) = −16

OpenStudy (anonymous):

−2(x − 1) = −16 2(x − 1) = 16 (x − 1) = 16/2 (x − 1) = 8 x=8+1 x=9

OpenStudy (anonymous):

Divide by -2 both the sides first and then add 1 to both the sides, you will get the answer..

OpenStudy (anonymous):

−2(x − 1) = −16 -2x+2=-16 -2x+18=0 x=9

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

Thanks you guys

OpenStudy (anonymous):

@mimosa101 did you understand??

OpenStudy (anonymous):

yes I did; thanks

OpenStudy (anonymous):

Welcome..